OpenStudy (anonymous):

Solve the equation below for x: Logx+log(x-2) = log3x Choices A. X =5 B. x = -2/3 C. X = 0,5 D. There are no solutions

OpenStudy (anonymous):

Remember log(xy) = log(x) + log(y). Try solving it now

OpenStudy (anonymous):

So would it be a? 5?

OpenStudy (anonymous):

Looks good to me, nice work : )

OpenStudy (anonymous):

No, thank you

OpenStudy (anonymous):

\[\log(x)(x-2) = \log3x\] Use his property to find out. \[\log(x^{2} -2x) = \log3x\]

OpenStudy (anonymous):

There is one more solution other than 5.

OpenStudy (anonymous):

But your right yes well done